OpenStudy (anonymous):

A family spends 20% of its monthly income on food, 23% on rent, and 42% on other expenses and saves the balance. If the family saves $360 per month, what is its monthly income?

OpenStudy (anonymous):

% thsy save = 100 - (20+23+42) = 15 % monthly income = (360 + 100) / 15 =

OpenStudy (anonymous):

sorry its ( 360 * 100) / 15 = 2400 dollars

OpenStudy (anonymous):

follow that ok ?